Murdock Industries uses a periodic inventory system and the LIFO retail method to estimate its ending inventories.

The following data has been summarized for December 31, 2018: Cost Retail Inventory, January 1 $ 113,000 $ 162,000 Purchases 355,534 538,500 Net markups 14,300 Net markdowns 8,500 Net sales 519,000 Required: Estimate the LIFO cost of ending inventory. Assume stable retail prices during the period. (Do not round your "Cost-to-retail percentage" and round your final answer to the nearest whole number.)

Estimated ending inventory:
Answer #1

In Last in First Out (LIFO) Method , the latest or more recent costs of products purchased (or produced) are the first costs expensed as the cost of goods sold. This means that the costs of the oldest products will be reported as inventory.

1.Cost of Opening Inventory : 162000/113000 = 1.4336

2.Cost of 355534 Purchases at 14300 Markup , Markup per unit = 14300/355534 = 0.0402

So , Cost of Purchase = 1.4336+0.0402 = 1.4738

3.Cost of 538500 Purchases at 8500 Markdown , Markdown Per Unit = 8500/538500 = 0.0158

So , Cost Of Purchase = 1.4336-0.0158 = 1.4178

Estimated Ending Inventory = {(538500-519000)*1.4178} + (355534*1.4738) +162000

=27647+523986+162000 = 713633

Estimate LIFO Cost is $ 713,633